# Genomic Next-Token Predictors are In-Context Learners

### v2

@@ 4.1 Main Results
Figure 2: Few-shot performance of Qwen3 and Evo2 models. (a) Evo2 model performance with respect to log(shots). All models monotonically improve – the 7B and 40B have roughly equivalent performance, and the 1B trails behind them. (b) Qwen3 model performance with respect to log(shots). All models improve, but not always monotonically. Smaller models struggle in 4-16 shot range. (c) At comparable sizes, Evo2 outperforms Qwen3. (d) Averaged performance across both model families shows consistent improvement with respect to log(shots). All models exceed the mode baseline shown in gray color.
@@
Evo2 outperforms Qwen3 at comparable sizes. Evo2 1B beats Qwen3 1.7B and Evo2 7B beats Qwen3 8B at higher shot counts (Fig. 2c). Averaging within families reinforces the same picture: comparable performance in the 1–4 shot range, Evo2 pulling ahead around 8–16 shots, and the gap persisting at higher shot counts (Fig. 2d.)
@@ 4.2 Analysis of ICL Capability Across Model Families and Sizes
As shown in Fig. 3, both Qwen3 and Evo2 exhibit clear performance gains with increasing scale. Qwen3 displays a strong positive correlation between accuracy and parameter count ($p<0.05$ for all shot counts $\geq$16, one-sided $t$-test on slope), while Evo2 shows a distinct jump from 1B to 7B but little improvement beyond that. This suggests that in-context program induction becomes more robust with scale, though saturation may occur once models reach sufficient capacity. A detailed statistical analysis is provided in Appendix D.
@@ 4.3 ICL Sensitivity to Task Complexity: BitLoad Analysis
Figure 4: Accuracy vs. BitLoad averaged across all tasks (BitLoad; Eq.3). Qwen declines sharply with increasing BitLoad, while Evo degrades more gradually, indicating greater robustness. Details in §4.3.
@@
where $x^{\oplus i}$ denotes $x$ with bit $i$ flipped, and $f_{j}(x)$ returns the $j$-th output bit. Intuitively, it counts the number of bit positions whose perturbation changes the output. So, the larger the BitLoad of a function, the harder it is since it requires the model to attend to more bits.
@@ 4.4 Qualitative Analysis of Functional and Behavioral Differences
Differential skill profiles. To sharpen contrasts, we identify the ten tasks most favoring each model. Qwen’s advantages are concentrated in simple shifts and aggregation tasks. It outperforms Evo by 37.5% on the "minority" operation (output all 1s if zeros >> ones), and by a similar margin on two parity-based tasks requiring counting the number of 1s. These trends suggest that Qwen may be better at reasoning over global properties of bitstrings. Its superiority on simple shifts may also be explained by the extreme rarity of frame shift mutations in DNA due to how catastrophic they are – a single nucleotide offset can decimate an entire protein. This is empirically supported by the fact that single-nucleotide deletions/shifts increase perplexity far more than other common mutations when presented to Evo2 (evo2report).
@@ 5 Discussion
    ICL’s emergence is due to data distributional properties: The distributional properties of data, such as “parallel structures” in human language pretraining data (chen2024parallel), its compositional structure (hahn2023theory), “burstiness” (chan2022data) and other such properties (wibisono2024context; reddy2023mechanistic) may be of importance (and perhaps necessary) for the emergence of ICL.
@@
    ICL’s emergence is due to a compression mechanism: The large-scale compression mechanism during massive pretraining might drive ICL (elmoznino2024complexity; elmoznino2024context; hahn2023theory).
@@
    ICL’s emergence may require specific architectural properties: While Transformers might be better suited for ICL than LSTMs (xie2021explanation), evidence is mixed (lee2023exploring), and non-Transformer models have also demonstrated ICL capabilities (grazzi2024mamba; park2024can).
@@
What are the implications of our findings on the frameworks to explain how ICL operates? We next consider the major perspectives that seek to explain how ICL operates. One view holds that ICL functions as a mix of task learning and task retrieval, with demonstrations serving either to recall pretrained capabilities or to enable learning on the fly (pan2023incontext; lin2024dual; wang2024investigating; fang2025iclciphersquantifyinglearning). Our symbolic reasoning tasks, instantiated in both linguistic and genomic domains, provide direct evidence for this task learning mode, aligning with this hypothesis and prior work (pan2023incontext; fang2025iclciphersquantifyinglearning). Because these tasks do not depend on pretrained semantic priors, they do not invoke task retrieval, offering limited insight into the Bayesian view that interprets ICL as implicit inference over latent concepts (xie2021explanation; panwar2023context; wang2024large; jiang2024context). Meanwhile, our results remain agnostic toward the optimization-based hypothesis, which posits that ICL implements an implicit gradient-descent-like process (akyurek2022learning; ahn2023transformers; mahankali2023one; li2023closeness), as well as the induction-based account, which attributes ICL to specialized “circuits” for performing inductive generalization (elhage2021mathematical; olsson2022context; wang2023label; bansal2023rethinking; ren2024identifying). Together, our results most strongly support the presence of genuine task learning within ICL.
Could Evo2’s tokenizer grant it a unique advantage on these tasks? Evo2’s absolute advantage over Qwen3 may be explained by the fact that Evo2’s entire pretraining distribution consists of long-context reasoning over a set of only a few tokens, whereas this kind of in-context operation is far more alien to Qwen3. Further analysis would be needed to verify if this is what grants Evo2 its absolute advantage. Ideally, a model family trained on a third modality with a vocabulary between a few nucleotides and hundreds of thousands of words could be employed.
@@ 6 Conclusion
Another promising direction is mechanistic interpretability: analyzing Evo2’s internal activations to identify which circuits enable ICL and how the model aggregates context to predict the next nucleotide. Comparing these mechanisms to known induction circuits (elhage2021mathematical) in LLMs could reveal whether analogous structures exist across architectures – a question that connects biological and linguistic ICL dynamics.
Finally, this work motivates searching for ICL in other non-linguistic modalities – time series (das2024decoderonlyfoundationmodeltimeseries), system logs (akhauri2025performancepredictionlargesystems), physics simulations (pmlr-v267-holzschuh25a), chess games (ruoss2024amortizedplanninglargescaletransformers), and climate projections (duncan2025samudracefastaccuratecoupled). Each offers a structured, patterned substrate that could support its own form of contextual reasoning. These diverse modalities, each with their unique structure and constraints, suggest a rich world of non-linguistic ICL capability waiting to be explored, and this work represents a maiden voyage into these extremely interesting waters.
@@ Appendix D Meta-Regression for ICL Efficacy with Number of Demonstrations
Figure 5: Few-shot behavior and scaling trends across Qwen3 and Evo2.
@@
Appendix E Analysis of ICL with Program Synthesis Tasks: Bit-Diversity
@@
Understandable Mistakes. A potential confound is what we will call "understandable mistakes". These mistakes occur when the model outputs an incorrect answer that would be correct for some other programs given the few-shot context. Formally, a model’s output $y$ (see 3.4 for notation):
$y=M\Big(e_{1}\to f(e_{1}),e_{2}\to f(e_{2}),\dots,e_{N}\to f(e_{N}),x\Big)$
is an understandable mistake if $y\neq f(x)$ but there exists $f_{2}\in F$ such that $\forall e_{i}\quad f(e_{i})=f_{2}(e_{i})$ and $y=f_{2}(x)$.

### v3

@@ 4.1 Main Results
Figure 2: Few-shot performance of Qwen3 and Evo2 models. (a) Evo2 model performance with respect to log(shots). All models monotonically improve – the 7B and 40B have roughly equivalent performance, and the 1B trails behind them. (b) Qwen3 model performance with respect to log(shots). All models improve, but not always monotonically. Smaller models struggle in 4-16 shot range. (c) At comparable sizes, Evo2 outperforms Qwen3. (d) Averaged performance across both model families shows consistent improvement with respect to log(shots). All models exceed the mode baseline shown in gray color. The exact accuracies and error bars (bootstrap-based standard errors) are included in Appendix C.
@@
Evo2 outperforms Qwen3 at comparable sizes. Evo2 1B beats Qwen3 1.7B and Evo2 7B beats Qwen3 8B at higher shot counts (Fig. 2c). Averaging within families reinforces the same picture: comparable performance in the 1–4 shot range, Evo2 pulling ahead around 8–16 shots, and the gap persisting at higher shot counts (Fig. 2d).
@@ 4.2 Analysis of ICL Capability Across Model Families and Sizes
As shown in Fig. 3, both Qwen3 and Evo2 exhibit clear performance gains with increasing scale. Qwen3 displays a strong positive correlation between accuracy and parameter count ($p<0.05$ for all shot counts $\geq$16, one-sided $t$-test on slope), while Evo2 shows a distinct jump from 1B to 7B but little improvement beyond that. This suggests that in-context program induction becomes more robust with scale, though saturation may occur once models reach sufficient capacity (A detailed statistical analysis in Appendix D).
@@ 4.3 ICL Sensitivity to Task Complexity: BitLoad Analysis
Figure 4: Accuracy vs. BitLoad averaged across all tasks (BitLoad; Eq.3). Qwen declines sharply, while Evo degrades more gradually. Details in §4.3.
@@
where $x^{\oplus i}$ denotes $x$ with bit $i$ flipped, and $f_{j}(x)$ returns the $j$-th output bit. Intuitively, it counts the number of bit positions whose perturbation changes the output. So, the larger the BitLoad of a function, the harder it is since it requires the model to attend to more bits. This metric is similar to the existing statistical measure of "relevant features", just defined specifically for our bitstring manipulation tasks. (Nevo and El-Yaniv, 2002).
@@ 4.4 Qualitative Analysis of Functional and Behavioral Differences
Differential skill profiles. To sharpen contrasts, we identify the ten tasks most favoring each model. Qwen’s advantages are concentrated in simple shifts and aggregation tasks. It outperforms Evo by 37.5% on the "minority" operation (output all 1s if zeros >> ones), and by a similar margin on two parity-based tasks requiring counting the number of 1s. These trends suggest that Qwen may be better at reasoning over global properties of bitstrings. Its superiority on simple shifts may also be explained by the extreme rarity of frame shift mutations in DNA due to how catastrophic they are – a single nucleotide offset can decimate an entire protein. This is empirically supported by the fact that single-nucleotide deletions/shifts increase perplexity far more than other common mutations when presented to Evo2 (Brixi et al., 2025).
@@ 5 Discussion
    ICL’s emergence is due to data distributional properties: The distributional properties of data, such as “parallel structures” in human language pretraining data (Chen et al., 2024), its compositional structure (Hahn and Goyal, 2023), “burstiness” (Chan et al., 2022) and other such properties (Wibisono and Wang, 2024; Reddy, 2023) may be of importance (and perhaps necessary) for the emergence of ICL.
@@
    ICL’s emergence is due to a compression mechanism: The large-scale compression mechanism during massive pretraining might drive ICL (Elmoznino et al., 2024a; b; Hahn and Goyal, 2023).
@@
    ICL’s emergence may require specific architectural properties: While Transformers might be better suited for ICL than LSTMs (Xie et al., 2021), evidence is mixed (Lee et al., 2023), and non-Transformer models have also demonstrated ICL capabilities (Grazzi et al., 2024; Park et al., 2024).
@@
What are the implications of our findings on the frameworks to explain how ICL operates? We next consider the major perspectives that seek to explain how ICL operates. One view holds that ICL functions as a mix of task learning and task retrieval, with demonstrations serving either to recall pretrained capabilities or to enable learning on the fly (Pan et al., 2023; Lin and Lee, 2024; Wang et al., 2024; Fang et al., 2025). Our symbolic reasoning tasks, instantiated in both linguistic and genomic domains, provide direct evidence for this task learning mode, aligning with this hypothesis and prior work (Pan et al., 2023; Fang et al., 2025). Because these tasks do not depend on pretrained semantic priors, they do not invoke task retrieval, offering limited insight into the Bayesian view that interprets ICL as implicit inference over latent concepts (Xie et al., 2021; Panwar et al., 2023; Wang et al., 2023b; Jiang et al., 2024). Meanwhile, our results remain agnostic toward the optimization-based hypothesis, which posits that ICL implements an implicit gradient-descent-like process (Akyürek et al., 2022; Ahn et al., 2023; Mahankali et al., 2023; Li et al., 2023b), as well as the induction-based account, which attributes ICL to specialized “circuits” for performing inductive generalization (Elhage et al., 2021; Olsson et al., 2022b; Wang et al., 2023a; Bansal et al., 2023; Ren et al., 2024). Together, our results most strongly support the presence of genuine task learning within ICL.
Does Evo2 have an innate advantage on these tasks? Possibly, for multiple reasons. First, though Evo2’s is trained on less tokens total, all of Evo2’s training tokens are long sequences of repeated nucleotides, and very little of Qwen3’s training tokens are long sequences of repeated digits. Second, Evo2’s StripedHyena2 architecture was found to to significantly outperform a vanilla transformer on long DNA sequences (Brixi et al., 2025). This could give Evo2 an innate advantage on long contexts containing the same few symbols vs Qwen3. Our result results do not imply Evo2’s ICL ability is superior to Qwen3’s, and we concede that our few-shot prompting setup may be biased toward Evo2. Attempts to make the task more legible to Qwen3, however, ran into the confound that Qwen3 has pretraining exposure to bitstring manipulation. Any prompting setup that included 0s and 1s immediately resulted in an extreme increase in Qwen3’s performance. Future work is needed to establish a method that controls for Evo2’s structural advantages without granting Qwen3 an unfair edge via its pretraining knowledge.
@@ 6 Conclusion
Another promising direction is mechanistic interpretability: analyzing Evo2’s internal activations to identify which circuits enable ICL and how the model aggregates context to predict the next nucleotide. Comparing these mechanisms to known induction circuits (Elhage et al., 2021) in LLMs could reveal whether analogous structures exist across architectures – a question that connects biological and linguistic ICL dynamics.
Finally, this work motivates searching for ICL in other non-linguistic modalities – time series (Das et al., 2024), system logs (Akhauri et al., 2025), physics simulations (Holzschuh et al., 2025), chess games (Ruoss et al., 2024), and climate projections (Duncan et al., 2025). Each offers a structured, patterned substrate that could support its own form of contextual reasoning. These diverse modalities, each with their unique structure and constraints, suggest a rich world of non-linguistic ICL capability waiting to be explored, and this work represents a maiden voyage into these extremely interesting waters.
@@
Acknowledgment
This work is supported by ONR grant (N0001424-1-2089). We acknowledge the use of computational resources on the DSAI cluster. We sincerely thank the JHU CLSP and DSAI communities for their helpful comments and feedback.
@@ Appendix D Meta-Regression for ICL Efficacy with Number of Demonstrations
Figure 5: Few-shot behavior and scaling trends across Qwen3 and Evo2.
Appendix E Demonstration of Evo2’s ICL Ability on a Genomic Task
We demonstrate that Evo2 7B can perform few-shot in-context learning (ICL) on human_nontata_promoters, a binary genomic classification task in which the model must predict whether a 251 nucleotide-long human DNA sequence is a promoter. To make the task non-trivial, sequences in the positive class are restricted to non-TATA promoters, removing the “TATA box”, a pattern that serves as an obvious, exploitable marker (Umarov and Solovyev, 2017; Grešová et al., 2023).
For each test query, we construct a few-shot prompt by concatenating balanced examples from the training set (equal numbers of promoters and non-promoters), followed by a held-out test sequence. Each training example is immediately followed by an encoded label. The label consists of a 24 nucleotide buffer (a single nucleotide repeated 24 times) and then a 24 nucleotide label run (another nucleotide repeated 24 times). The buffer nucleotide and the two label nucleotides (for class 0 vs. class 1) are re-sampled uniformly at random without replacement from $\{A,C,G,T\}$ for every trial, so the mapping changes across evaluations. This prevents any nucleotide-specific biasing effects.
To predict the test label, we score two candidate prompts - one where the test sequence is followed by the class-0 label and one followed by the class-1 label - and choose the label with lower perplexity on the label run (i.e., computed only over the final 24 label nucleotides, excluding the buffer). This lets us measure whether Evo2 can infer the label mapping from the few-shot context and apply it to the query sequence.
We observe monotonic improvement with respect to shot count, with accuracy rising from below random baseline at 1-shot (48.2%) to 63.6% at 16-shot. These gains further rise to 68.6% after 256 shots. A saturated power law achieves an $R^{2}$ of 0.998 in the 2-256 shot regime, which is characteristic of few-shot ICL. Furthermore, we show via ablation that the improvement is due to correctly-labeled examples presented in context—randomly permuting the labels in the 16-shot regime drops performance from 63.6% back to random chance (50.2%).
Figure 6: Performance of Evo2 7B on human_nontata_promoters at varying shot counts. Error bars are $\pm$ one standard error. Evo2 7B’s performance shows clear monotonic improvements with respect to shot count.
Appendix F Analysis of ICL with Program Synthesis Tasks: Bit-Diversity
@@
Understandable Mistakes. A potential confound is what we will call "understandable mistakes". These mistakes occur when the model outputs an incorrect answer that would be correct for some other programs given the few-shot context. Formally, a model’s output $y$ (see 3.4 for notation): $y=M\Big(e_{1}\to f(e_{1}),e_{2}\to f(e_{2}),\dots,e_{N}\to f(e_{N}),x\Big)$ is an understandable mistake if $y\neq f(x)$ but there exists $f_{2}\in F$ such that $\forall e_{i}\quad f(e_{i})=f_{2}(e_{i})$ and $y=f_{2}(x)$.